OpenStudy (anonymous):

Are 20, 25 and 30 an example of a Pythagorean triplet? Yes or No, Explain your answer and show work.

OpenStudy (amistre64):

does the sum of the smaller squares equal the larger square?

OpenStudy (amistre64):

also, scaling them down might be prdent if doing it by hand

OpenStudy (amistre64):

20, 25 and 30 -> 4, 5, and 6 work the smaller set

OpenStudy (anonymous):

No, the given numbers are not Pythagorean Triplets. following is the reason \[30^2=900------(1)\] \[20^2 +25^2=400 + 625 = 1025----(2)\] From eq (1) and (2) it is clear that \[30^2 \neq 20^2 + 25^2\] Thus given numbers are not Pythagorean triplet.